find the polynomial equation that has the following roots: 0, -1, and 5
Answer by stanbon(75887)

find the polynomial equation that has the following roots: 0, -1, and 5
---------------------
Rule: If r is a root, x-r is a factor.
------
Your Problem:
f(x) = x(x+1)(x-5)
----
f(x) = x(x^2-4x-5)
----
f(x) = x^3-4x^2-5
========================
Cheers,
Stan H.
================